### v1

# An Optimism-based Approach to Online Evaluation of Generative Models

Abstract
Existing frameworks for evaluating and comparing generative models typically target an offline setting, where the evaluator has access to full batches of data produced by the models. However, in many practical scenarios, the goal is to identify the best model using the fewest generated samples to minimize the costs of querying data from the models. Such an online comparison is challenging with current offline assessment methods. In this work, we propose an online evaluation framework to find the generative model that maximizes a standard assessment score among a group of available models. Our method uses an optimism-based multi-armed bandit framework to identify the model producing data with the highest evaluation score, quantifying the quality and diversity of generated data. Specifically, we study the online assessment of generative models based on the Fréchet Inception Distance (FID) and Inception Score (IS) metrics and propose the FID-UCB and IS-UCB algorithms leveraging the upper confidence bound approach in online learning. We prove sub-linear regret bounds for these algorithms and present numerical results on standard image datasets, demonstrating their effectiveness in identifying the score-maximizing generative model.
1 Introduction
Deep generative models have achieved astonishing results across a wide array of machine learning datasets. Quantitative comparisons between generative models, trained using different methods and architectures, are commonly performed by evaluating assessment metrics such as Fréchet Inception Distance (FID) [17] and Inception Score (IS) [37]. Due to the growing applications of generative models to various learning tasks, the machine learning community has continuously adapted evaluation methodologies to better suit the characteristics of the newly introduced applications.
A common characteristic of standard evaluation frameworks for deep generative models is their offline assessment process, which requires a full batch of generated data for assigning scores to the models. While this offline evaluation does not incur significant costs for moderate-sized generative models, producing large batches of samples from large-scale models can be costly. In particular, generating a large batch of high-resolution image or video data could be expensive and hinder the application of existing evaluation scores for ranking generative models.
In this work, we focus on the online evaluation of generative modeling schemes, where we consider a group of generative models and attempt to identify the model with the best score by assessing the fewest number of produced data. By limiting the number of generated samples in the assessment process, online evaluation can save on the costs associated with producing large batches of samples. Additionally, online evaluation can significantly reduce the time and computational expenses required to identify a well-performing model in assessing a large group of generative models.
To measure the performance of an online evaluation algorithm, we use the regret notion widely-used in the online learning literature. In the online learning task, the evaluation algorithm selects a generative model and observes a mini-batch of samples generated by that model in each round. Our goal is to minimize the algorithm’s regret, defined as the cumulative difference between the scores of the selected models and the best possible score from the model set. Therefore, the learner’s regret quantifies the cost of generating samples from sub-optimal models. Our goal is to develop online learning algorithms that result in lower regret values in the assessment of generated data.
To address the described online evaluation of generative models, we utilize the multi-armed bandit framework. In a standard multi-armed bandit problem, the online learner seeks to identify the arm with the highest expected value of a random score. While this formulation has been widely considered in many learning settings, it cannot solve the online evaluation of generative models, as standard IS and FID metrics do not simplify to the expectation of a random variable and represent a non-linear function of the data distribution.
By deriving concentration bounds for the FID and IS scores, we propose the optimism-based FID-UCB and IS-UCB algorithms to address the online assessment of generative models. These algorithms apply the upper confidence bound (UCB) approach using data-dependent bounds that we establish for the FID and IS scores. Furthermore, we analytically bound the regret of the IS-UCB and FID-UCB methods, demonstrating their sub-linear regret growth assuming a full-rank covariance matrix of the embedded real data.
We discuss the results of several numerical applications of FID-UCB and IS-UCB to standard image datasets and generative modeling frameworks. We compare the performance of these algorithms with the greedy algorithm, which selects the model with the highest estimated score, and a naive-UCB baseline that applies the UCB algorithm with a data-independent upper confidence bound. Our numerical results show a significant improvement in our proposed algorithms compared to the greedy and naive-UCB baselines. Additionally, we analyze the performance of FID-UCB with different image data embeddings, including InceptionNet.V3 [40], CLIP [7], and DINOv2 [33]. Our empirical results indicate satisfactory performance under these standard embeddings. The following summarizes the main contributions of this work:
  • Proposing an online evaluation framework for generative models that aims to minimize the regret of misidentifying the score-maximizing model from online generated data.
  • Developing the FID-UCB and IS-UCB algorithms by applying the upper-confidence-bound framework to our data-dependent estimation of the scores.
  • Proving sub-linear regret bounds for the proposed FID-UCB and IS-UCB algorithms.
  • Demonstrating satisfactory empirical performance of FID-UCB and IS-UCB in comparison to the greedy and naive-UCB baseline algorithms.
2 Related Work
Assessment of deep generative models. The evaluation of generative models has been extensively studied in the literature. Several evaluation metrics are proposed, including distance-based metrics such as Wasserstein critic [3], Fréchet Inception Distance (FID) [17] and Kernel Inception Distance (KID) [5], and diversity/quality-based metrics such as Precision/Recall [36, 27], density and coverage [32], VENDI [13], and RKE [20]. In addition, the related works develop metrics quantifying the generalizability of the generative models, including the authenticity score [2], the FLD score [21], the Rarity score [16], and the KEN score [45] measuring the novelty of the generated samples. In this paper, we primarily focus on FID and Inception score, which have been frequently used for evaluating generative models.
Role of embeddings in the quantitative evaluation results. Due to the high-dimensionality of images, evaluation of the generated images mostly relies on the embeddings extracted by pretrained networks on the ImageNet dataset, e.g., InceptionNet.V3 [40]. However, [32] shows that such pretrained-embeddings could exhibit unexpected behaviors. Recently, several large pretrained models have been proposed, including DINOv2 [33] and CLIP [7]. [39] shows that DINOv2-ViT-L/14 enables more interpretable evaluation of generative models. In addition, [28] demonstrates that FID scores computed based on the embedding extracted by CLIP agree more with human-based assessments. In this paper, we provide the numerical results for the mentioned embeddings extracted by different pretrained models.
Online Learning using diversity-related evaluation metrics. Online learning is a sequential decision-making framework where an agent aims to minimize a cumulative loss function revealed to her sequentially. One popular setting is the multi-armed bandit (MAB), whose study dates back to the work of [29, 41]. At each step, the agent chooses among several arms, each associated with a reward distribution, and aims to maximize a pre-specified performance metric. The primary concern of this body of literature considers maximizing the expected return [1, 4, 6]. Recent works study performance metrics related to the variance or entropy of the reward distribution, including the mean-variance criterion [38, 47] in risk-sensitive MAB, and informational MAB (IMAB) where the agent maximizes the entropy rewards [44]. However, to the best of our knowledge, the evaluation of generative models has not been exclusively studied in an online learning context.
Online training of generative models. Several related works focus on training generative adversarial networks (GANs) [14] using online learning frameworks. [15] proposes to train semi-shallow GANs using the Follow-the-Regularized-Leader (FTRL) approach. [10] shows that optimistic mirror decent (OMD) can be applied to address the limit cycling problem in training Wasserstein GANs (WGANs). Also, the recent paper [34] studies the no-regret behaviors of large language model (LLM) agents. This reference proposes an unsupervised training loss, whose minimization could automatically result in known no-regret learning algorithms. On the other hand, our focus is on the evaluation of generative models which does not concern the models’ training.
3 Preliminaries
3.1 Inception Score
Inception score (IS) is a standard metric for evaluating generative models, defined as
$\textup{IS}:=\exp\left\{\mathbb{E}_{X_{\textup{g}}\sim p_{\textup{g}}}[\textup{KL}(p_{Y|X_{\textup{g}}}\|p_{Y_{\textup{g}}})]\right\},$ (1)
where $X_{\textup{g}}\sim p_{\textup{g}}$ is a generated image, $p_{Y|X_{\textup{g}}}$ is the conditional class distribution assigned by the InceptionNet.V3 pretrained on ImageNet [40], and $p_{Y_{\textup{g}}}=\mathbb{E}_{X_{\textup{g}}\sim p_{\textup{g}}}[p_{Y|X_{\textup{g}}}]$ is the marginal class distribution. Further, we have that $\textup{IS}=\exp[I(X_{\textup{g}};Y_{\textup{g}})]=\exp[H(Y_{\textup{g}})-H(Y_{\textup{g}}|X_{\textup{g}})]$, where $I(\cdot;\cdot)$ is the mutual information, and $H(\cdot)$ is the Shannon entropy. A higher IS implies that the images generated by the model have higher diversity, since $p_{Y_{\textup{g}}}$ would be more uniformly distributed to increase $H(Y_{\textup{g}})$, and possess higher fidelity, because $p_{Y|X_{\textup{g}}}$ is closer to a one-hot vector to enforce a smaller $H(Y_{\textup{g}}|X_{\textup{g}})$.
3.2 Fréchet Inception Distance
Fréchet Inception Distance (FID) is another standard metric for evaluating generative models. Let $f(x)\in{\mathbb{R}}^{d}$ denote the $d$-dimensional feature of an image $x$, which is extracted by the InceptionNet.V3 embedding. FID is the Fréchet distance [12] between the feature distributions of the generated images $p_{f(X_{\textup{g}})}$ and the real images $p_{f(X_{\textup{r}})}$, defined as
$\textup{FID}=\|\mu_{\textup{g}}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\Sigma_{\textup{g}}+\Sigma_{\textup{r}}-2(\Sigma_{\textup{g}}\Sigma_{\textup{r}})^{\frac{1}{2}}]$ (2)
assuming that $p_{f(X_{\textup{g}})}={\mathcal{N}}(\mu_{\textup{g}},\Sigma_{\textup{g}})$ and $p_{f(X_{\textup{r}})}={\mathcal{N}}(\mu_{\textup{r}},\Sigma_{\textup{r}})$ are Gaussian distributions.
4 Online Evaluation of Generative Models
In this section, we introduce the framework of online evaluation of generative models, which is given in Protocol 1. We denote by ${\mathcal{G}}:=[G]$ the set of generative models. For each generator $g\in[G]$, we denote by $p_{g}\in\Delta({\mathcal{X}})$ its generative distribution over the space ${\mathcal{X}}$, which can be texts or images. Given an evaluation metric, e.g., FID or IS, the corresponding score of the generator $g\in[G]$ is denoted by $\nu_{g}$. The evaluation proceeds in $T$ steps. At each step $t\in[T]$, the evaluating algorithm ${\mathcal{A}}$ picks a generator $g_{t}\in[G]$ and collects a batch of generated samples $\{x_{t}^{j}\sim p_{g}\}_{j=1}^{b}$, where $b\in{\mathbb{N}}_{+}$ is the (fixed) batch size. The evaluator aims to minimize the regret
$\textup{Regret}(T)=\sum_{t=1}^{T}(\nu^{*}-\nu_{g_{t}}),$ (3)
where $\nu^{*}:=\operatorname*{arg\,max}_{g\in[G]}\nu_{g}$ (if the higher the score the better).
Regarding the challenges of online evaluation of generative models, note that the empirical estimation of the score could be biased and generator-dependent. In addition, the generative distribution typically lies in a high-dimensional space, which makes it difficult to estimate the score from limited data. Moreover, the evaluation metrics often incorporate higher-order information of the generative distribution (e.g., computing FID requires the covariance matrix). Hence, analyzing the concentration properties of the score estimation would be challenging.
Protocol 1 Online Evaluation of Generative Models
[…]
1:  Initialize: estimated score $\widehat{\nu}_{g}$ and generated samples ${\mathcal{H}}_{g}\leftarrow\varnothing$ for each $g\in[G]$.
2:  for step $t=1,2,\cdots,T$  do
3:     Pick generator $g_{t}\sim{\mathcal{A}}$.
4:     Generate a batch $\{x_{t}^{j}\sim p_{g_{t}}\}_{j=1}^{b}$ of samples from $g_{t}$.
5:     Update ${\mathcal{H}}_{g_{t}}\leftarrow{\mathcal{H}}_{g_{t}}\cup\{x_{t}^{j}\}_{j=1}^{b}$ and estimated score $\widehat{\nu}_{g_{t}}$.
6:  end for
5 Online Evaluation based on Fréchet Inception Distance
In this section, we consider the online evaluation of generative models by Fréchet Inception Distance (FID). Given $n\in{\mathbb{N}}_{+}$ generated images $x^{1},\cdots,x^{n}\sim p_{g}$ queried from model $g$, the empirical FID is computed by
$\widetilde{\textup{FID}}_{g}^{n}=\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\widehat{\Sigma}_{g}^{n}+\Sigma_{\textup{r}}-2(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}],$ (4)
where
$\widehat{\Sigma}_{g}^{n}=\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\widehat{\mu}_{g}^{n})(f(x^{i})-\widehat{\mu}_{g}^{n})^{\top},\qquad\widehat{\mu}_{g}^{n}=\frac{1}{n}\sum_{i=1}^{n}f(x^{i})$ (5)
are the empirical covariance matrix and the mean vector, respectively, and $f(x^{i})\in{\mathbb{R}}^{d}$ is the feature of the $i$-th generated image which is extracted by, e.g., the InceptionNet.V3. It has been shown that the empirical FID (4) is biased differently depending on the generator, which makes it difficult to compare different generators for a fixed sample size [9]. Hence, the FID-based evaluation is typically performed for a large batch of generated samples (e.g., $10\sim 50$k) to reduce the effect of the bias, which can be sample-inefficient and costly for a large number of generators.
To enable sample-efficient online evaluation of generative models, we adapt the optimism-based online learning framework to the FID score. To this end, we first derive an optimistic FID score in the following theorem. We defer the theorem’s proof to Appendix A.1.
Theorem 5.1 (Generator-dependent optimistic FID score).
Assume the covariance matrix $\Sigma_{\textup{r}}\succ\bm{0}$ is positive definite. Then, with probability at least $1-\delta$, we have
$\widehat{\textup{FID}}_{g}^{n}=\widetilde{\textup{FID}}_{g}^{n}-{\mathcal{B}}^{n}_{g}\leq\textup{FID}_{g}.$ (6)
Here, the bonus is given by
${\mathcal{B}}_{g}^{n}:=$ $C^{n}_{g}\cdot\sqrt{\frac{2{\mathcal{I}}_{g}}{n}}\left(32\log\left(\frac{24d}{\delta}\right)\right)^{\frac{1}{4}}+\textup{Tr}[\Sigma_{g}]\sqrt{\frac{8}{n}\log\left(\frac{6d}{\delta}\right)}$ (7)
$+\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\left(20\sigma_{g}^{2}\|\Sigma_{g}\|_{2}\sqrt{\frac{4\bm{r}(\Sigma_{g})+\log(3/\delta)}{n}}+\frac{2{\mathcal{I}}_{g}}{n}\sqrt{32\log\left(\frac{24d}{\delta}\right)}\right)},$
where $C_{g}^{n}:=2(\|\mu_{g}-\mu_{\textup{r}}\|_{2}+\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\frac{1}{n}\sum_{i=1}^{n}\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2})$, ${\mathcal{I}}_{g}:=\sum_{j=1}^{d}\sum_{l=1}^{d}|\Sigma_{g}[l,j]|$, and $\bm{r}(\Sigma_{g}):=\frac{\textup{Tr}[\Sigma_{g}]}{\|\Sigma_{g}\|_{2}}$.
Algorithm 1 FID-UCB for Generative Model Evaluation
[…]
1:  Initialize: $\widehat{\textup{FID}}_{g}\leftarrow-\infty,{\mathcal{H}}_{g}\leftarrow\varnothing,n_{g}\leftarrow 0$ for each $g\in[G]$.
2:  for step $t=1,2,\cdots,T$ do
3:     Pick model $g_{t}\leftarrow\operatorname*{arg\,min}_{g\in{\mathcal{G}}}\widehat{\textup{FID}}_{g}$.
4:     Model $g_{t}$ generates batch samples $\{x_{t}^{j}\in{\mathbb{R}}^{D}\sim g_{t}\}_{j=1}^{b}$.
5:     Update ${\mathcal{H}}_{g_{t}}\leftarrow{\mathcal{H}}_{g_{t}}\cup\{f(x_{t}^{j})\}_{j=1}^{b}$ and visitation $n_{g_{t}}\leftarrow n_{g_{t}}+b$.
6:     Update $\widehat{\mu}_{g_{t}}$ and $\widehat{\Sigma}_{g_{t}}$.
7:     Compute the optimistic FID score
$\widehat{\textup{FID}}_{g_{t}}\leftarrow\|\widehat{\mu}_{g_{t}}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}\left[\widehat{\Sigma}_{g_{t}}+\Sigma_{\textup{r}}-2(\widehat{\Sigma}_{g_{t}}\Sigma_{\textup{r}})^{\frac{1}{2}}\right]-{\mathcal{B}}_{g_{t}}^{n_{g_{t}}}.$
8:  end for
We remark that the bonus term (7) is: 1) generator-dependent through parameters including ${\mathcal{I}}_{g}$, $\sigma_{g}^{2}$, $\textup{Tr}[\Sigma_{g}]$, and $\|\Sigma_{g}\|_{2}$, 2) dimension-free in the sense that the dimension $d$ only appears in the logarithmic term, which facilitates a sample-efficient evaluation. In practice, these parameters can be estimated from the queried data.
Based on Theorem 5.1, we propose FID-UCB in Algorithm 1 as an FID-based online evaluation algorithm. At the beginning, the estimated FID scores are initialized to be negative infinity (line 1), and hence each generator is explored for at least one time. The evaluation proceeds iterative, where at each iteration $t\in[T]$, the evaluator picks model $g_{t}$ with the lowest estimated FID and queries a batch of images from the model (lines $3\sim 4$). Then, the estimated FID of generator $g_{t}$ is updated (lines $5\sim 7$). Next, we show that FID-UCB attains sub-linear regret bound. We defer the theorem’s proof to Appendix A.2.
Theorem 5.2 (Regret of FID-UCB).
Under the same conditions in Theorem A.1, with probability at least $1-\delta$, the regret of the UCB-FID algorithm after $T$ steps is bounded by
$\widetilde{O}\left(\max_{g}\left\{G\cdot{\mathcal{I}}_{g}\cdot\log(T),\sqrt{G({\mathcal{I}}_{g}+\textup{Tr}[\Sigma_{g}])T},(\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\cdot\sigma_{g}\|\Sigma_{g}\|_{2}^{\frac{1}{2}}\bm{r}^{\frac{1}{4}}(\Sigma_{g}))\cdot G^{\frac{1}{4}}T^{\frac{3}{4}}\right\}\right),$ (8)
where logarithmic factors are hidden in the notation $\widetilde{O}(\cdot)$.
6 Online Evaluation based on Inception Score
In this section, we focus on evaluating generative models by Inception score (IS), which is given by $\textup{IS}=\exp[I(X_{g};Y_{g})]$, where $X_{g}\sim p_{g}$ is the generated image, and $Y_{g}$ is the class assigned by the InceptionNet.V3. Given $n\in{\mathbb{N}}_{+}$ generated images $x^{1},\cdots,x^{n}\sim p_{g}$ queried from a generator $g$, the empirical IS is computed by
$\widetilde{\textup{IS}}_{g}^{n}=\exp\left\{\widehat{H}^{n}(Y_{g})-\widehat{H}^{n}(Y_{g}|X_{g})\right\},$ (9)
where
$\widehat{H}^{n}(Y_{g})=H\left(\frac{1}{n}\sum_{i=1}^{n}p_{Y|x^{i}}\right),\ \widehat{H}^{n}(Y_{g}|X_{g})=\frac{1}{n}\sum_{i=1}^{n}H(Y|x^{i}),$ (10)
are the empirical entropy of the marginal $d$-class distribution $p_{Y_{g}}$ and the conditional $d$-class distribution $p_{Y|X_{g}}$ ($d=1000$ for InceptionNet.V3).
For any $j\in[d]$, let $\widehat{V}^{n}(p_{Y|X_{g}}[j]):={\mathbb{V}}(p_{Y|x^{1}}[j],\cdots,p_{Y|x^{n}}[j])$ denote the sample variance for the probability that the generated sample is assigned to the $j$-th class. To derive an optimistic IS, we first define the optimistic marginal class distribution denoted by
$\widehat{p}_{Y_{g}}^{n}:=\textup{Clip}_{e^{-1}}(\widetilde{p}^{n}_{Y_{g}},\bm{\epsilon}^{n}_{g})$ (11)
where $\bm{\epsilon}^{n}_{g}\in{\mathbb{R}}_{+}^{d}$ is a $d$-dimensional vector whose $j$-th element is given by
$\bm{\epsilon}^{n}_{g}[j]:=\sqrt{\frac{2\widehat{V}^{n}(p_{Y|X_{g}}[j])}{n}\log\left(\frac{4d}{\delta}\right)}+\frac{7}{3(n-1)}\log\left(\frac{4d}{\delta}\right),$ (12)
and $\textup{Clip}_{e^{-1}}(p,\bm{\epsilon})$ is the following element-wise operator
$\begin{cases}p[j]+\frac{e^{-1}-p[j]}{|e^{-1}-p[j]|}\bm{\epsilon}[j]&\textup{, if }|e^{-1}-p[j]|\geq\bm{\epsilon}[j]\\ e^{-1}&\textup{, otherwise.}\end{cases}$ (13)
The optimistic marginal class distribution ensures that ${\mathcal{E}}(\widehat{p}_{Y_{g}}^{n}):=-\sum_{j}\widehat{p}_{Y_{g}}^{n}[j]\cdot\log(\widehat{p}_{Y_{g}}^{n}[j])\geq H(Y_{g})$ with high probability (see Lemma B.1 in the Appendix B.3). Next, we derive a generator-dependent optimistic IS in the following theorem. We defer the theorem’s proof to Appendix B.1.
Theorem 6.1 (Generator-dependent Optimistic IS).
Let
$\widehat{\textup{IS}}^{n}_{g}:=\exp\left\{{\mathcal{E}}\left(\widehat{p}_{Y_{g}}^{n}\right)-\widehat{H}^{n}(Y_{g}|X_{g})+\sqrt{\frac{2\widehat{V}^{n}(H(Y_{g}|X_{g}))}{n}\log\left(\frac{4}{\delta}\right)}+\frac{7C^{\prime}}{3(n-1)}\log\left(\frac{4}{\delta}\right)\right\},$ (14)
where $\widehat{H}^{n}(Y_{g}|X_{g})$ is defined in Equation (10), $\widehat{p}_{Y_{g}}^{n}$ the optimistic marginal class distribution (11), parameter $C^{\prime}$ satisfies that $P(|H(Y_{g}|X_{g})|\leq C^{\prime})=1$, and $\widehat{V}^{n}(H(Y_{g}|X_{g}))={\mathbb{V}}(H(Y|x^{1}),\cdots,H(Y|x^{n}))$ is the empirical variance for the entropy of the conditional class distribution. Then, with probability at least $1-\delta$, we have that $\widehat{\textup{IS}}^{n}_{g}\geq\textup{IS}_{g}$.
[…]
1:  Initialize: $\widehat{\textup{IS}}_{g}\leftarrow\infty,\widehat{H}(Y_{g}|X_{g})\leftarrow 0,\widetilde{p}_{Y_{g}}\leftarrow\bm{0}^{d},n_{g}\leftarrow 0$ for each $g\in[G]$.
2:  for step $t=1,2,\cdots,T$ do
3:     Pick model $g_{t}\leftarrow\operatorname*{arg\,max}_{g\in{\mathcal{G}}}\widehat{\textup{IS}}_{g}$.
4:     Model $g_{t}$ generates batch samples $\{x_{t}^{j}\sim p_{g_{t}}\}_{j=1}^{b}$.
5:     Update
$\widetilde{p}_{Y_{g_{t}}}\leftarrow\frac{N_{g}\cdot\widetilde{p}_{Y_{g}}+\sum_{j=1}^{b}p_{Y|x_{t}^{i}}}{n_{g_{t}}+b},\ \widehat{H}(Y_{g_{t}}|X_{g_{t}})\leftarrow\frac{N_{g}\cdot\widehat{H}(Y_{g_{t}}|X_{g_{t}})+\sum_{j=1}^{b}H(Y|x_{t}^{i})}{n_{g_{t}}+b}.$
6:     Update visitation $N_{g_{t}}\leftarrow N_{g_{t}}+b$.
7:     Update the empirical variances $\widehat{V}(H(Y_{g_{t}}|X_{g_{t}}))$ and $\widehat{V}(p_{Y|X_{g_{t}}}[j])$ for any $j\in[d]$.
8:     Compute the estimated IS according to Equation (14).
9:  end for
Algorithm 2 IS-UCB for Generative Model Evaluation
Based on Theorem 6.1, we propose IS-UCB in Algorithm 2, an optimism-based online IS evaluation algorithm. In the following theorem, we derive the regret bound of the IS-UCB algorithm, which shows that IS-UCB attains $\widetilde{O}(\sqrt{T})$ regret. We present the theorem’s proof in Appendix B.2.
Theorem 6.2 (Regret of IS-UCB).
With probability at least $1-\delta$, the regret of the IS-UCB algorithm after $T$ steps is bounded by
$\textup{Regret}(T)\leq\widetilde{O}\left(e^{C}\cdot\max\left\{dm,\left(\frac{d\log d}{4}\right)^{2}\right\}\sqrt{T}+dG\Delta^{-2}\log d\right),$ (15)
where $C:=O(de^{-1}+C^{\prime})$ is the upper bound of $\widehat{\textup{IS}}_{g}^{n}$, $m:=\max_{g\in[G],j\in[d]:p_{Y_{g}}[j]>0}\{|u^{\prime}(p_{Y_{g}}[j])|\}$, $\Delta:=\min_{g\in[G],j\in[d]:p_{Y_{g}}[j]\neq e^{-1}}\{|p_{Y_{g}}[j]-e^{-1}|\}$, and $u(x)=-x\log x$.
7 Experimental Results
Baseline methods. For both FID-based and IS-based evaluation, we compare the performance of our proposed FID-UCB and IS-UCB with the following two baselines, naive-UCB and the greedy algorithm. Naive-UCB is a simplification of FID-UCB and IS-UCB which is based on data-independent dimension-based concentration bounds, which do not exploit the generated data in evaluating the confidence bound. We present the exact estimators in Appendix D. On the other hand, the greedy algorithm always picks the generator with the lowest empirical FID (4) or the highest empirical IS (9). For all these algorithms, we initialize the estimated FID to be $\infty$- ∞ and the estimated IS to be +$\infty$+ ∞, and hence each generator will be explored at least once.
[…]
Experimental setup and performance metrics. We report three performance metrics at each step $t\in[T]$: 1) total regret, i.e., $\textup{Regret}(t)$, 2) average regret, i.e., $(1/t)\cdot\textup{Regret}(t)$, and 3) optimal pick ratio, i.e., the ratio $(1/t)\cdot\sum_{i=1}^{t}\bm{1}(g_{t}=g^{*})$ of picking the optimal model, which has the lowest empirical FID score or the highest empirical IS score for 50k generated images. For all the experiments, we use a batch size of 5, and the total evaluation step is $T=1,000$. Hence the total generated samples for each trial is $5$k. Results are averaged over 20 trials. Due to the limited space, we postpone parts of the results to Appendix C.
7.1 Empirical Results of Online FID-based Evaluation
Performance on pretrained generators under different embeddings. The results are summarized in Figure 1, where each row and each column correspond to one dataset and one performance metric, respectively. The results show that FID-UCB (blue) outperforms naive-UCB (orange) and the greedy algorithm (green). We observe that naive-UCB fails to distinguish between different generators and makes uniform selections on the CIFAR10 and the ImageNet datasets, while FID-UCB recognizes the optimal one. This shows that the generator-dependent optimistic FID (6) can better exploit the properties of the generator, which is key to the sample-efficient online evaluation. On the FFHQ dataset, all the generators attain very similar FID scores (the largest gap is less than 0.7), which may explain why FID-UCB has a similar performance to naive-UCB. We reran these experiments using DINOv2-ViT-L/14-based [33] and CLIP-based [7] embeddings. (The results are summarized in Figures 5 and 6 in Appendix C, respective.) The numerical results show that FID-UCB can maintain satisfactory performance over different embeddings.
CIFAR10 (G=5)
ImageNet (G=5)
[…]
Figure 1: Performance of FID-UCB, naive-UCB, and the greedy algorithm on various datasets. The embeddings of the images are extracted by the pretrained InceptionNet.V3.
Performance on variance-controlled generators. We summarize the results in Figure 2. FID-UCB (blue) significantly outperforms naive-UCB and the greedy algorithm. The results for DINOv2-based and CLIP-based FID are provided in Figures 5 and 6 in Appendix C. We ran the evaluation on AFHQ Dog dataset, and the results are presented in Figure 7 in the Appendix. Furthermore, we display the generated images to visualize the quality and diversity of the samples queried by different algorithms. The results suggest that FID-UCB consistently queries images that are diverse and of high quality. In contrast, naive-UCB kept querying a proportion of images from the collapsed generator (the one with the smallest truncation parameter), and the greedy algorithm even stuck to it.
FFHQ (G=20)
Figure 2: Performance of FID-UCB, naive-UCB, and the greedy algorithm for evaluating the variance-controlled generators on the FFHQ dataset.
Step 500
[…]
7.2 Results of Online IS-Based Evaluation
Performance on the pretrained generators. The results are summarized in Figure 3. IS-UCB (blue) attains significantly better performance than naive-UCB (orange) and the greedy algorithm (green) on CIFAR10 and FFHQ datasets.
CIFAR10 (G=5)
Figure 3: Performance of IS-UCB, naive-UCB, and the greedy algorithm on various datasets.
ImageNet (G=5) [Uncaptioned image]
[…]
Performance on variance-controlled generators. The results are summarized in Figure 4. Generated samples are plotted in Figure 8 in the Appendix. The results on the AFHQ Dog dataset are presented in Figure 9 in the Appendix.
FFHQ (G=20)
Figure 4: Performance of IS-UCB, naive-UCB, and the greedy algorithm for evaluating the variance-controlled generators on the FFHQ dataset.
8 Conclusion
In this work, we studied an online learning problem aiming to identify the generative model with the best evaluation score among a set of models. We proposed the optimism-based FID-UCB and IS-UCB algorithms to address the online learning task and showed satisfactory numerical results for their application to image-based generative models. An interesting future direction to our work is to explore the application of other online learning frameworks, such as Thompson sampling, to address the online evaluation problem. Also, while we mostly focused on standard image datasets for testing the proposed algorithms, the application of the proposed FID-UCB and IS-UCB to text-based and video-based generative models will be interesting for future studies. Such an application only requires applying text and video-based embeddings, as in Video IS [35], FVD [42], and FTD [19].
Appendix A Proofs in Section 5: FID Evaluation
A.1 Proof of Theorem 5.1: Optimistic FID Score
Proof.
The proof of Theorem 5.1 is based on Theorem A.1 and Lemma A.2 in the Appendix. Specifically, Inequality (16) in Theorem A.1 in Appendix A.3 decomposes the estimation error of the empirical FID score (4) into the errors in estimating the mean (i.e., $\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}$) and the covariance matrix (i.e., $\|\widehat{\Sigma}_{g}^{n}-\Sigma_{g}\|_{2}$). Next, Lemma A.2 in Appendix A.4 derives generator-dependent concentration errors for the mean and the covariance matrix. Combining ∎
A.2 Proof of Theorem 5.2: Regret of FID-UCB
Proof.
Recall that $g_{t}$ denotes the generator picked at the $t$-th step. For convenience, we denote by $\widehat{\textup{FID}}_{g_{t}}$ the optimistic FID of generator $g_{t}$ computed at step $t$. We denote by $n_{t}$ the number of images generated by model $g_{t}$ after $t$ steps. First, by Theorem A.2 and a union bound over $T$ steps, with probability at least $1-\delta$, we have that $\textup{FID}_{g}\geq\widehat{\textup{FID}}_{g}$ for any step $t\in[T]$ and $g\in[G]$. Hence, we have that
$\textup{Regret}(T)=\sum_{t=1}^{T}\left(\textup{FID}_{g_{t}}-\textup{FID}^{*}\right)\leq\sum_{t=1}^{T}\left(\textup{FID}_{g_{t}}-\widehat{\textup{FID}}_{g_{t}}\right).$
By the definition of $\widehat{\textup{FID}}_{g_{t}}$, we further derive that
$\textup{Regret}(T)\leq$ $\sum_{t=1}^{T}\left(\textup{FID}_{g_{t}}-\widetilde{\textup{FID}}_{g_{t}}^{n_{t}}+{\mathcal{B}}^{n_{t}}_{g_{t}}\right)\leq 2\sum_{t=1}^{T}{\mathcal{B}}^{n_{t}}_{g_{t}},$
where ${\mathcal{B}}^{n}_{g}$ is the bonus function given by Equation (7). Therefore, the regret is further bounded by
$\widetilde{O}\left(\max_{g}\left\{G\cdot{\mathcal{I}}_{g}\cdot\log(T),\sqrt{G({\mathcal{I}}_{g}+\textup{Tr}[\Sigma_{g}])\cdot T},(\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\cdot\sigma_{g}\|\Sigma_{g}\|_{2}^{\frac{1}{2}}\bm{r}^{\frac{1}{4}}(\Sigma_{g}))\cdot G^{\frac{1}{4}}T^{\frac{3}{4}}\right\}\right),$
where logarithmic factors are hidden in the notation $\widetilde{O}(\cdot)$. Here, we use the fact that
$\sum_{t=1}^{T}\frac{1}{n_{t}^{\alpha}}\leq\frac{(2b)^{\alpha}}{1-\alpha}\sum_{g=1}^{G}(n_{g}(T))^{1-\alpha}\leq\frac{(2b)^{\alpha}}{1-\alpha}G^{\alpha}T^{1-\alpha}\text{, where }0\leq\alpha<1,$
$\sum_{t=1}^{T}\frac{1}{n_{t}}\leq 2b\sum_{g=1}^{G}\log(n_{g}(T))\leq 2bG\log(T),$
where $b$ is the batch size, and $n_{g}(T)$ is the number of images generated by model $g$ at the last step $T$. ∎
A.3 Concentration of Empirical FID (4)
Theorem A.1 (Concentration of empirical FID (4)).
Assume the covariance matrix $\Sigma_{\textup{r}}\succ\bm{0}$ is positive strictly definite, and we denote by $\Sigma_{\textup{r}}=Q_{\textup{r}}^{\top}\Lambda_{\textup{r}}Q_{\textup{r}}$ its eigendecomposition. Then, with probability at least $1-\frac{\delta}{3}$, we have that
$\left|\widetilde{\textup{FID}}_{g}^{n}-\textup{FID}_{g}\right|\leq$ $\underbrace{(\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}+2\|\mu_{g}-\mu_{\textup{r}}\|_{2})\cdot\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}}_{\text{(1) error of $\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}$}}+\underbrace{\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}}_{\text{(2) error of $2\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$}}$ (16)
$+\underbrace{\textup{Tr}(\Sigma_{g})\sqrt{\frac{8}{n}\log\left(\frac{6d}{\delta}\right)}+\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}\cdot\left(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\frac{2}{n}\sum_{i=1}^{n}\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}\right)}_{\text{(3) error of $\textup{Tr}[\widehat{\Sigma}_{g}^{n}]$}}.$
Proof.
Recall that for any generator $g$, we have that
$\textup{FID}_{g}=\|\mu_{g}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\Sigma_{g}+\Sigma_{\textup{r}}-2(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}],$
$\widetilde{\textup{FID}}_{g}^{n}=\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}+\textup{Tr}[\widehat{\Sigma}_{g}^{n}+\Sigma_{\textup{r}}-2(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}].$
(1) Bound $\|\mu_{g}-\mu_{\textup{r}}\|_{2}^{2}-\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}$.  Note that
$\left|\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}-\|\mu_{g}-\mu_{\textup{r}}\|_{2}\right|\leq\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}.$
We derive
$\left|\|\mu_{g}-\mu_{\textup{r}}\|_{2}^{2}-\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}\right|$ (17)
$\leq$ $(\|\mu_{g}-\mu_{\textup{r}}\|_{2}+\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2})\cdot\left|\|\mu_{g}-\mu_{\textup{r}}\|_{2}-\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}\right|$
$\leq$ $(2\|\mu_{g}-\mu_{\textup{r}}\|_{2}+\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2})\cdot\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}.$
(2) Bound $\textup{Tr}[(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}-(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$.  By Theorem A.3 in Appendix A.5, if the covariance matrix $\Sigma_{\textup{r}}$ is positive strictly definite, then it holds that
$\left|\textup{Tr}[(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}]-\textup{Tr}\left[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}\right]\right|\leq\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{2\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}.$ (18)
(3) Bound $\textup{Tr}[\Sigma_{g}-\widehat{\Sigma}_{g}^{n}]$.  Note that
$\left|\textup{Tr}[\Sigma_{g}-\widehat{\Sigma}_{g}^{n}]\right|$
$=$ $\left|\mathbb{E}[\textup{Tr}[(f(X_{g})-\mu_{g})(f(X_{g})-\mu_{g})^{\top}]]-\frac{1}{n}\sum_{i=1}^{n}\textup{Tr}[(f(x^{i})-\widehat{\mu}_{g}^{n})((f(x^{i})-\widehat{\mu}_{g}^{n}))^{\top}]\right|$
$=$ $\left|\mathbb{E}[\|f(X_{g})-\mu_{g}\|_{2}^{2}]-\frac{1}{n}\sum_{i=1}^{n}\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}^{2}\right|$
$\leq$ $\left|\mathbb{E}[\|f(X_{g})-\mu_{g}\|_{2}^{2}]-\frac{1}{n}\sum_{i=1}^{n}\|f(x^{i})-\mu_{g}\|_{2}^{2}\right|+\left|\frac{1}{n}\sum_{i=1}^{n}(\|f(x^{i})-\mu_{g}\|_{2}^{2}-\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}^{2})\right|.$ (19)
Next, we bound the two terms in the last inequality. Let $Z_{g}:=f(X_{g})-\mu_{g}\sim{\mathcal{N}}(0,\Sigma_{g})$. For the first term, we have that
$\|f(X_{g})-\mu_{g}\|_{2}^{2}=\|Z_{g}\|_{2}^{2}=(\Sigma_{g}^{-\frac{1}{2}}Z_{g})^{\top}\Sigma_{g}(\Sigma_{g}^{-\frac{1}{2}}Z_{g})=(Q_{g}\Sigma_{g}^{-\frac{1}{2}}Z_{g})^{\top}\Lambda_{g}(Q_{g}\Sigma_{g}^{-\frac{1}{2}}Z_{g}),$
where we denote by $\Sigma_{g}:=Q_{g}^{\top}\Lambda_{g}Q_{g}$ the eigendecomposition of $\Sigma_{g}$. Note that the random vector $Q_{g}\Sigma_{g}^{-\frac{1}{2}}Z_{g}$ has distribution ${\mathcal{N}}(0,I)$. We have that
$\|f(X_{g})-\mu_{g}\|_{2}^{2}=\sum_{j=1}^{d}\lambda_{j}U_{j}^{2},$
where $\lambda_{j}$’s are the eigenvalues of $\Sigma_{g}$, $U_{i}\sim{\mathcal{N}}(0,1)$ are independent standard normal, and hence $U_{i}^{2}$’s are independent Chi-square random variables. Therefore, we have that
$\left|\mathbb{E}[\|f(X_{g})-\mu_{g}\|_{2}^{2}]-\frac{1}{n}\sum_{i=1}^{n}\|f(x^{i})-\mu_{g}\|_{2}^{2}\right|$
$=$ $\left|\sum_{j=1}^{d}\lambda_{j}\left(\frac{1}{n}\sum_{i=1}^{n}(U_{j}^{i})^{2}-\mathbb{E}[U_{j}^{2}]\right)\right|$
$\leq$ $\left(\sum_{j=1}^{d}\lambda_{j}\right)\cdot\max_{j=1,\cdots,d}\left|\frac{1}{n}\sum_{i=1}^{n}(U_{j}^{i})^{2}-1\right|$
$\leq$ $\textup{Tr}(\Sigma_{g})\sqrt{\frac{8}{n}\log\left(\frac{6d}{\delta}\right)},$ (20)
where the last inequality holds with probability at least $1-\frac{\delta}{3}$ by Lemma E.5. In addition, for the second term, we derive that
$\left|\frac{1}{n}\sum_{i=1}^{n}(\|f(x^{i})-\mu_{g}\|_{2}^{2}-\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}^{2})\right|$
$=$ $\left|\frac{1}{n}\sum_{i=1}^{n}(\|f(x^{i})-\mu_{g}\|_{2}+\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2})(\|f(x^{i})-\mu_{g}\|_{2}-\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2})\right|$
$\leq$ $\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}\cdot\left(\frac{1}{n}\sum_{i=1}^{n}(\|f(x^{i})-\mu_{g}\|_{2}+\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2})\right)$
$\leq$ $\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}\cdot\left(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\frac{2}{n}\sum_{i=1}^{n}\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}\right).$ (21)
Putting all together.  Therefore, combining Inequalities (17), (18), and (19), with probability at least $1-\frac{\delta}{3}$, it holds that
$\left|\widetilde{\textup{FID}}_{g}^{n}-\textup{FID}_{g}\right|\leq$ $\underbrace{(\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}+2\|\mu_{g}-\mu_{\textup{r}}\|_{2})\cdot\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}}_{\text{(1) error of $\|\widehat{\mu}_{g}^{n}-\mu_{\textup{r}}\|_{2}^{2}$}}+\underbrace{\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}}_{\text{(2) error of $2\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$}}$
$+\underbrace{\textup{Tr}(\Sigma_{g})\sqrt{\frac{8}{n}\log\left(\frac{6d}{\delta}\right)}+\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}\cdot\left(\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}+\frac{2}{n}\sum_{i=1}^{n}\|f(x^{i})-\widehat{\mu}_{g}^{n}\|_{2}\right)}_{\text{(3) error of $\textup{Tr}[\widehat{\Sigma}_{g}^{n}]$}},$
which concludes the proof. ∎
A.4 Concentration of Mean Vector and Covariance Matrix
Lemma A.2 (Concentration of mean vector and covariance matrix).
Under the same conditions in Theorem A.1, with probability at least $1-\frac{2\delta}{3}$, we have that
$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}\leq\sqrt{\frac{2{\mathcal{I}}_{g}}{n}}\left(32\log\left(\frac{24d}{\delta}\right)\right)^{\frac{1}{4}},\text{ where }{\mathcal{I}}_{g}:=\sum_{j=1}^{d}\sum_{l=1}^{d}|\Sigma_{g}[l,j]|$ (22)
and
$\left\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\right\|_{2}\leq 20\sigma_{g}^{2}\|\Sigma_{g}\|_{2}\sqrt{\frac{4\bm{r}(\Sigma_{g})+\log(3/\delta)}{n}}+\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2},$ (23)
where $\sigma_{g}^{2}:=\max_{i\in[d]}\Sigma_{g}[i,i]$ is the largest variance for the elements in $f(X_{g})$, and $\bm{r}(\Sigma_{g})=\frac{\textup{Tr}[\Sigma_{g}]}{\|\Sigma_{g}\|_{2}}$ is the effective rank of $\Sigma_{g}$.
Proof.
1. Concentration of mean vector.  Note that $\Sigma_{g}^{\frac{1}{2}}Z$ where $Z\sim{\mathcal{N}}(0,I)$ has a multivariate ${\mathcal{N}}(0,\Sigma_{g})$ distribution, which is the same as $f(X_{g})-\mu_{g}$. We have that
$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}=\left\|\frac{1}{n}\sum_{i=1}^{n}(\Sigma_{g}^{\frac{1}{2}}Z_{i})\right\|_{2}=\left\|\frac{1}{n}\sum_{i=1}^{n}Z_{i}\right\|_{\Sigma_{g}},$
where $\|y\|_{\Sigma_{g}}:=\sqrt{\langle y,\Sigma_{g}y\rangle}$ is the matrix norm. Let $\widehat{Z}:=\frac{1}{n}\sum_{i=1}^{n}Z_{i}$ denote the mean vector. We further obtain that
$\left\|\widehat{Z}\right\|_{\Sigma_{g}}^{2}=\textup{Tr}\left((\widehat{Z})^{\top}\Sigma_{g}\widehat{Z}\right)=\textup{Tr}\left(\Sigma_{g}\widehat{Z}(\widehat{Z})^{\top}\right)=\textup{Tr}(\Sigma_{g}D)+\textup{Tr}(\Sigma_{g}M),$ (24)
where $D:=\textup{diag}(\widehat{Z}(\widehat{Z})^{\top})$ and $M=\widehat{Z}(\widehat{Z})^{\top}-D$. For the first term, note that each diagonal element has the same distribution $\frac{1}{n^{2}}(\sum_{i=1}^{n}z_{i})^{2}$, where $z_{1},\cdots,z_{n}\sim{\mathcal{N}}(0,1)$ are i.i.d standard Gaussian. Then, with probability at least $1-\frac{\delta}{6}$, we have that
$\textup{Tr}(\Sigma_{g}D)\leq\|D\|_{2}\textup{Tr}(\Sigma_{g})\leq\textup{Tr}(\Sigma_{g})\left(\frac{1}{n}\sqrt{32\log\left(\frac{12d}{\delta}\right)}\right),$ (25)
where the first inequality holds by that $\Sigma_{g}$ is PSD, and the second inequality holds by Lemma E.10. For the second term,
$\textup{Tr}(\Sigma_{g}M)=$ $\sum_{j=1}^{d}\sum_{l=1,l\neq j}^{d}\Sigma_{g}[l,j]\cdot M_{j,l}$
$=$ $\sum_{j=1}^{d}\sum_{l=1,l\neq j}^{d}\Sigma_{g}[l,j]\cdot\left(\frac{\sum_{i=1}^{n}z_{j,i}}{n}\right)\left(\frac{\sum_{i=1}^{n}z_{l,i}}{n}\right)$
$\leq$ $\sum_{j=1}^{d}\sum_{l=1,l\neq j}^{d}|\Sigma_{g}[l,j]|\cdot\max_{j,l}\left|\left(\frac{\sum_{i=1}^{n}z_{j,i}}{n}\right)\left(\frac{\sum_{i=1}^{n}z_{l,i}}{n}\right)\right|$
where $z_{j,1},\cdots,z_{j,n},z_{l,1},\cdots,z_{l,n}\sim{\mathcal{N}}(0,1)$ are i.i.d standard Gaussian. By Lemma E.11, with probability at least $1-\frac{\delta}{6}$, it holds that
$\textup{Tr}(\Sigma_{g}M)\leq\left(\sum_{j=1}^{d}\sum_{l=1,l\neq j}^{d}|\Sigma_{g}[l,j]|\right)\frac{1}{n}\sqrt{8\log\left(\frac{24d}{\delta}\right)}$ (26)
Combining Inequalities (25) and (26), we conclude that with probability at least $1-\frac{\delta}{3}$, it holds that
$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|^{2}_{2}\leq\left(\sum_{j=1}^{d}\sum_{l=1}^{d}|\Sigma_{g}[l,j]|\right)\left(\frac{2}{n}\sqrt{32\log\left(\frac{24d}{\delta}\right)}\right)$ (27)
2. Concentration of sample covariance matrix. we have that
$\left\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\right\|_{2}$
$=$ $\left\|\Sigma_{g}-\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\widehat{\mu}_{g}^{n})(f(x^{i})-\widehat{\mu}_{g}^{n})^{\top}\right\|_{2}$
$=$ $\left\|\Sigma_{g}-\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\mu_{g}+\mu_{g}-\widehat{\mu}_{g}^{n})(f(x^{i})-\mu_{g}+\mu_{g}-\widehat{\mu}_{g}^{n})^{\top}\right\|_{2}$
$\leq$ $\left\|\Sigma_{g}-\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\mu_{g})(f(x^{i})-\mu_{g})^{\top}\right\|_{2}+\left\|(\mu_{g}-\widehat{\mu}_{g}^{n})(\mu_{g}-\widehat{\mu}_{g}^{n})^{\top}\right\|_{2},$ (28)
where the last inequality holds by the triangle inequality of the L2 norm of the matrix. Note that the zero-mean random vector $f(X_{g})-\mu_{g}$ has a multivariate ${\mathcal{N}}(0,\Sigma_{g})$ distribution (under the assumption of FID). For the first term, we evoke Lemma E.7. Then, with probability at least $1-\frac{\delta}{3}$, it holds that
$\left\|\Sigma_{g}-\frac{1}{n}\sum_{i=1}^{n}(f(x^{i})-\mu_{g})(f(x^{i})-\mu_{g})^{\top}\right\|_{2}\leq 20\sigma_{g}^{2}\|\Sigma_{g}\|_{2}\sqrt{\frac{4\bm{r}(\Sigma_{g})+\log(3/\delta)}{n}}.$ (29)
For the second term, we have that
$\left\|(\mu_{g}-\widehat{\mu}_{g}^{n})(\mu_{g}-\widehat{\mu}_{g}^{n})^{\top}\right\|_{2}=\|\mu_{g}-\widehat{\mu}_{g}^{n}\|_{2}^{2},$
which concludes the proof. ∎
A.5 Dimension-free Concentration of $\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$
Theorem A.3 (Dimension-free concentration of $\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$).
Under the same conditions in Theorem A.1, we have that
$\left|\textup{Tr}[(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}]-\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]\right|\leq\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{2\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}}.$ (30)
Proof.
Let $\eta=\|\Sigma_{g}-\widehat{\Sigma}_{g}^{n}\|_{2}$. We have that
$\textup{Tr}[(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}]-\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]$
$\leq$ $\textup{Tr}\left[\sqrt{B^{\top}(\Sigma_{g}-\widehat{\Sigma}_{g}^{n}+\widehat{\Sigma}_{g}^{n}+\eta I)B}\right]-\textup{Tr}\left[\sqrt{B^{\top}\widehat{\Sigma}_{g}^{n}B}\right]$
$\leq$ $\textup{Tr}\left[\sqrt{B^{\top}(\Sigma_{g}-\widehat{\Sigma}_{g}^{n}+\eta I)B}+\sqrt{B^{\top}\widehat{\Sigma}_{g}^{n}B}\right]-\textup{Tr}\left[\sqrt{B^{\top}\widehat{\Sigma}_{g}^{n}B}\right]$
$\leq$ $\textup{Tr}\left[\sqrt{2\eta B^{\top}B}\right]=\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{2\eta},$
where the first inequality holds by Lemma E.10, the second inequality holds by Lemma E.11 and the fact that both matrices $B^{\top}(\Sigma_{g}-\widehat{\Sigma}_{g}^{n}+\eta I)B\succeq\bm{0},B^{\top}\widehat{\Sigma}_{g}^{n}B\succeq\bm{0}$ are PSD, and the last inequality holds by Lemma E.10 and $B^{\top}(2\eta I)B\succeq B^{\top}(\Sigma_{g}-\widehat{\Sigma}_{g}^{n}+\eta I)B$. By the same analysis, we can also derive that
$\textup{Tr}[(\widehat{\Sigma}_{g}^{n}\Sigma_{\textup{r}})^{\frac{1}{2}}]-\textup{Tr}[(\Sigma_{g}\Sigma_{\textup{r}})^{\frac{1}{2}}]\leq\textup{Tr}\left[\sqrt{2\eta B^{\top}B}\right]=\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{2\eta},$
which concludes the proof. ∎
Appendix B Proofs in Section 6: IS Evaluation
B.1 Proof of Theorem 6.1: Generator-Dependent Optimistic IS
Proof.
By Lemma B.2, we have that with probability at least $1-\frac{\delta}{2}$, ${\mathcal{E}}(\widehat{p}_{Y_{g}}^{n})\geq H(Y_{g})$. To show Theorem 6.1, it suffices to show that
$H_{g}(Y_{g}|X_{g})\geq\widehat{H}^{n}(Y_{g}|X_{g})-\sqrt{\frac{2\widehat{V}_{g}^{n}(H(Y_{g}|X_{g}))}{n}\log\left(\frac{4}{\delta}\right)}-\frac{7C^{\prime}}{3(n-1)}\log\left(\frac{4}{\delta}\right)$
with probability at least $1-\frac{\delta}{2}$, which holds by Theorem E.4. Therefore, we conclude the proof. ∎
B.2 Proof of Theorem 6.2: Regret of IS-UCB
Proof.
By Theorem 6.1 and union bound over $T$ steps, with probability at least $1-\delta$, it holds that $\widehat{\textup{IS}}_{g}\geq\textup{IS}_{g}$ for all steps $t\in[T]$ and $g\in[G]$, where $\widehat{\textup{IS}}_{g}$ is given by Equation (14). For convenience, we denote by $\widehat{\textup{IS}}_{g_{t}}$ the optimistic IS of generator $g_{t}$ computed at the $t$-th step. Hence, we have that
$\textup{IS}^{*}-\textup{IS}_{g_{t}}\leq\widehat{\textup{IS}}_{g_{t}}-\textup{IS}_{g_{t}}=e^{{\mathcal{E}}\left(\widehat{p}_{Y_{g_{t}}}\right)-\widehat{H}(Y_{g_{t}}|X_{g_{t}})+{\mathcal{B}}_{g_{t}}}-\textup{IS}_{g_{t}},$
where
${\mathcal{B}}_{g_{t}}:=\sqrt{\frac{2\widehat{V}_{g}^{n}(H(Y_{g}|X_{g}))}{n}\log\left(\frac{4}{\delta}\right)}+\frac{7C^{\prime}}{3(n-1)}\log\left(\frac{4}{\delta}\right).$
Let ${\mathcal{T}}_{1},{\mathcal{T}}_{2}\subset[T]$ be two disjoint set of steps such that ${\mathcal{T}}_{2}=[T]\backslash{\mathcal{T}}_{1}$. Specifically, ${\mathcal{T}}_{1}$ contains steps where no element of the empirical marginal class distribution $\widetilde{p}_{Y_{g_{t}}}$ is clipped to $e^{-1}$, i.e., $|e^{-1}-\widetilde{p}_{Y_{g_{t}}}[j]|\geq\bm{\epsilon}_{g_{t}}[j]$ for all $[j]\in d$. Let $\textup{IS}^{*}=\operatorname*{arg\,max}_{g\in[G]}\textup{IS}_{g}$ denote the optimal Inception score. Hence, we have that
$\textup{Regret}(T)=\sum_{t\in{\mathcal{T}}_{1}}(\textup{IS}^{*}-\textup{IS}_{g_{t}})+\sum_{t\in{\mathcal{T}}_{2}}(\textup{IS}^{*}-\textup{IS}_{g_{t}}).$
Recall $u(x)=-x\log x$ for $x\in{\mathbb{R}}_{+}$. For the first part, we further derive that
$\sum_{t\in{\mathcal{T}}_{1}}(\textup{IS}^{*}-\textup{IS}_{g_{t}})$
$\leq$ $\sum_{t\in{\mathcal{T}}_{1}}\left(\exp\{{\mathcal{E}}\left(\widehat{p}_{Y_{g_{t}}}\right)-\widehat{H}(Y_{g_{t}}|X_{g_{t}})+{\mathcal{B}}_{g_{t}}\}-\exp\{H(Y_{g})-H(Y_{g}|X_{g})\}\right)$
$\leq$ $e^{C}\cdot\sum_{t\in{\mathcal{T}}_{1}}\left({\mathcal{E}}\left(\widehat{p}_{Y_{g_{t}}}\right)-\widehat{H}(Y_{g_{t}}|X_{g_{t}})+{\mathcal{B}}_{g_{t}}-H(Y_{g})+H(Y_{g}|X_{g})\right)$
$\leq$ $e^{C}\cdot\sum_{t\in{\mathcal{T}}_{1}}\left(\sum_{j=1}^{d}\left(u(\widehat{p}_{Y_{g_{t}}}[j])-u(p_{Y_{g_{t}}}[j])\right)+2{\mathcal{B}}_{g_{t}}\right)$
$\leq$ $O\left(e^{C}\cdot\sum_{t\in{\mathcal{T}}_{1}}\left(m\cdot\sum_{j=1}^{d}|\widehat{p}_{Y_{g_{t}}}[j]-p_{Y_{g_{t}}}[j]|+{\mathcal{B}}_{g_{t}}\right)\right)$
$\leq$ $O\left(e^{C}\cdot\sum_{t\in{\mathcal{T}}_{1}}\left(m\cdot\sum_{j=1}^{d}\bm{\epsilon}_{g_{t}}[j]+{\mathcal{B}}_{g_{t}}\right)\right)$
$\leq$ $\widetilde{O}\left(e^{C}\cdot\sum_{t=1}^{T}\left(\max\left\{dm,\frac{d\log d}{2}\right\}\sqrt{\frac{1}{N_{g_{t}}}}+\frac{C^{\prime}}{N_{g_{t}}}\right)\right),$ (31)
where $m=\max_{g\in[G],j\in[d]:p_{Y_{g}}[j]>0}\{|u^{\prime}(p_{Y_{g}}[j])|\}$, and $C=O(de^{-1}+C^{\prime})$ is the upper bound of ${\mathcal{E}}(\widehat{p}_{Y_{g_{t}}})-\widehat{H}(Y_{g_{t}}|X_{g_{t}})+{\mathcal{B}}_{g_{t}}$. For the second part, note that the number of steps where at least one element of $\widetilde{p}_{Y_{g_{t}}}$ is clipped to at most $\widetilde{O}(G\Delta^{-2})$, where $\Delta=\min_{g\in[G],j\in[d]:p_{Y_{g}}[j]\neq e^{-1}}\{|p_{Y_{g}}[j]-e^{-1}|\}$. Therefore, we have that
$\textup{Regret}(T)\leq\widetilde{O}\left(dG\Delta^{-2}\log d+e^{C}\cdot\max\left\{dm,\left(\frac{d\log d}{4}\right)^{2}\right\}\sqrt{T}\right),$
which concludes the proof. ∎
B.3 Optimistic marginal class distribution
Theorem B.1 (Optimistic marginal class distribution).
Let $\widetilde{p}^{n}_{Y_{g}}:=\frac{1}{n}\sum_{i=1}^{n}p_{Y_{g}|x^{i}}$ denote the empirical marginal class distribution. Let
$\widehat{p}_{Y_{g}}^{n}=\textup{Clip}_{e^{-1}}\left(\widetilde{p}^{n}_{Y_{g}},|\widetilde{p}^{n}_{Y_{g}}-p_{Y_{g}}|\right),$
where $\textup{Clip}_{e^{-1}}(p,\bm{\epsilon})$ is the following element-wise operator
$\begin{cases}p[j]+\frac{e^{-1}-p[j]}{|e^{-1}-p[j]|}\bm{\epsilon}[j]&\textup{, if }|e^{-1}-p[j]|\geq\bm{\epsilon}[j]\\ e^{-1}&\textup{, otherwise.}\end{cases}$
for any vectors $p,\bm{\epsilon}\in{\mathbb{R}}^{d}_{+}$. Then, we have that ${\mathcal{E}}(\widehat{p}_{Y_{g}}^{n})\geq H(Y_{g})$.
Proof.
Note that the function $u(z)=-z\log z$ is concave and attains maximum at $z=e^{-1}$. It suffices to show that $u(\widehat{p}_{Y_{g}}[j])\geq u(p_{Y_{g}}[j])$ for all $j\in[d]$. If $|e^{-1}-\widehat{p}_{Y_{g}}^{n}[j]|<|\widetilde{p}^{n}_{Y_{g}}[j]-p_{Y_{g}}[j]|$, then $\widehat{p}_{Y_{g}}[j]=e^{-1}$, which ensures that $u(\widehat{p}_{Y_{g}}[j])\geq u(p_{Y_{g}}[j])$. In addition, if $|e^{-1}-\widehat{p}_{Y_{g}}^{n}[j]|\geq|\widetilde{p}^{n}_{Y_{g}}[j]-p_{Y_{g}}[j]|$, then
$\widehat{p}_{Y_{g}}^{n}[j]=\widetilde{p}^{n}_{Y_{g}}[j]+\frac{e^{-1}-\widetilde{p}^{n}_{Y_{g}}[j]}{|e^{-1}-\widetilde{p}^{n}_{Y_{g}}[j]|}|\widetilde{p}^{n}_{Y_{g}}[j]-p_{Y_{g}}[j]|=\begin{cases}\widetilde{p}^{n}_{Y_{g}}[j]+|\widetilde{p}^{n}_{Y_{g}}[j]-p_{Y_{g}}[j]|,&\textup{ if }p_{Y_{g}}[j]<e^{-1}\\ \widetilde{p}^{n}_{Y_{g}}[j]-|\widetilde{p}^{n}_{Y_{g}}[j]-p_{Y_{g}}[j]|,&\textup{ otherwise.}\end{cases}.$
The first case ensures that $p_{Y_{g}}[j]<\widehat{p}_{Y_{g}}^{n}[j]\leq e^{-1}$, and the second case ensures that $p_{Y_{g}}[j]>\widehat{p}_{Y_{g}}^{n}[j]\geq e^{-1}$. Both cases satisfy that $u(\widehat{p}_{Y_{g}}^{n}[j])\geq u(p_{Y_{g}}[j])$. Therefore, we have that $H(Y_{g})={\mathcal{E}}(p_{Y_{g}})=\sum_{j=1}^{d}u(p_{Y_{g}}[j])\leq\sum_{j=1}^{d}u(\widehat{p}_{Y_{g}}^{n}[j])={\mathcal{E}}(\widehat{p}_{Y_{g}}^{n})$, which concludes the proof. ∎
B.4 Data-dependent optimistic marginal class distribution
Lemma B.2 (Optimistic marginal class distribution).
Let $x^{1},\cdots,x^{n}\sim p_{g}$ be $n$ generated images from generator $g$. Define $\bm{\epsilon}^{n}_{g}\in{\mathbb{R}}_{+}^{d}$ denote the error vector whose $j$-th element is given by
$\bm{\epsilon}^{n}_{g}[j]=\sqrt{\frac{2\widehat{V}^{n}(p_{Y|X_{g}}[j])}{n}\log\left(\frac{4d}{\delta}\right)}+\frac{7}{3(n-1)}\log\left(\frac{4d}{\delta}\right),$
where $\widehat{V}^{n}(p_{Y|X_{g}}[j])={\mathbb{V}}(p_{Y|x^{1}}[j],\cdots,p_{Y|x^{n}}[j])$ is the empirical variance for the $j$-th class density. Then, with probability at least $1-\frac{\delta}{2}$, we have that
$\widehat{p}_{Y_{g}}^{n}=\textup{Clip}_{e^{-1}}\left(\widetilde{p}^{n}_{Y_{g}},\bm{\epsilon}^{n}_{g}\right)$
satisfies that ${\mathcal{E}}(\widehat{p}_{Y_{g}}^{n})\geq H(Y_{g})$.
Proof.
It suffices that show that $\bm{\epsilon}_{g}^{n}[j]\geq|\widehat{p}_{Y_{g}}^{n}[j]-p_{Y_{g}}[j]|$ with probability at least $1-\frac{\delta}{2}$. We evoke Theorem E.4 and conclude the proof. ∎
Appendix C Additional Experimental Results
C.1 Results of FID-Based Online Evaluation
1. Embeddings Extracted by DINOv2 [33].
We run FID-UCB, naive UCB, and the greedy algorithm where the feature of an image is extracted by DINOv2-ViT-L/14. The pretrained weights are downloaded from the official repository (licensed under Apache License 2.0). [^3] https://github.com/facebookresearch/dinov2/tree/main The results are summarized in Figure 5.
CIFAR10 (G=5)
ImageNet (G=5)
FFHQ (G=5)
[…]
Figure 5: Performance of FID-UCB, naive UCB, and the greedy algorithm on various datasets, where the last row corresponds to the evaluation of the variance-controlled generators. The embeddings of the images are extracted by the pretrained DINOv2-ViT-L/14 model.
2. Embeddings Extracted by CLIP [7].
We run FID-UCB, naive UCB, and the greedy algorithm where the feature of an image is extracted by CLIP. The pretrained weights are downloaded from the official repository. [^4] https://github.com/mlfoundations/open_clip. Copyright (c) 2012-2021 Gabriel Ilharco, Mitchell Wortsman, Nicholas Carlini, Rohan Taori, Achal Dave, Vaishaal Shankar, John Miller, Hongseok Namkoong, Hannaneh Hajishirzi, Ali Farhadi, Ludwig Schmidt. The results are summarized in Figure 6. We note that in the FFHQ dataset, FID-UCB attains a low optimal pick ratio as the optimal generator and the second-best one attains almost the same CLIP-based FID.
CIFAR10 (G=5)
ImageNet (G=5)
FFHQ (G=5)
[…]
Figure 6: Performance of FID-UCB, naive UCB, and the greedy algorithm on various datasets, where the last row corresponds to the evaluation of the variance-controlled generators. The embeddings of images are extracted by the pretrained CLIP model.
3. Results on AFHQ Dog dataset.
The results are summarized in Figure 7.
AFHQ Dog (G=20)
Figure 7: Performance of FID-UCB, naive UCB, and the greedy algorithm for evaluating the variance-controlled generators on AFHQ Dog dataset. The embeddings of images are extracted by InceptionNet.V3.
C.2 Results of IS-Based Online Evaluation
1. Diversity and quality of the generated samples.
We examine the diversity and quality of the generated samples queried by IS-UCB, naive UCB, and the greedy algorithm (i.e., experiment in Figure 4). The results are summarized in Figure 8.
Step 500
[…]
Figure 8: Generated images queried by IS-UCB, naive UCB, and the greedy algorithm.
2. Results on AFHQ Dog dataset.
The results are summarized in Figure 9.
AFHQ Dog (G=20)
Figure 9: Performance of IS-UCB, naive UCB, and the greedy algorithm for evaluating the variance-controlled generators on AFHQ dataset.
Appendix D Additional Experimental details
1. Additional experiment setups.
All the experiments are run on a single Nvidia GeForce RTX 3090 GPU. For FID-based evaluation, each trial takes at most 2 hours to finish, depending on the dataset and the embedding pretrained model. For IS-based evaluation, each trial takes less than 30 minutes to finish.
2. List of pretrained generative models/generated image data.
For the CIFAR10 dataset, we compare pretrained models including SNGAN, ACGAN, BigGAN, ContraGAN, and BigGAN-CR. For the ImageNet dataset, we compare pretrained models including StyleGAN3, SNGAN, BigGAN, ContraGAN, and ReACGAN. The pretrained weights are downloaded from the StudioGAN repository [22]. For the FFHQ dataset, we compare StyleSwin, StyleNAT, StyleGAN2-ADA, StyleGAN-XL, and LDM, where we utilize the generated image datasets downloaded from the dgm-eval repository [39].
3. Implementation details of naive UCB for FID-based evaluation.
In FID-based evaluation, naive UCB upper bounds the error in estimating the mean vector by
$\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}=\sqrt{\sum_{j=1}^{d}(\widehat{\mu}_{g}^{n}[j]-\mu_{g}[j])^{2}}\leq\sigma_{g}\sqrt{\frac{d}{n}\log\left(\frac{2d}{\delta}\right)}$
which holds by the assumption of FID score that $f(X_{g})\sim{\mathcal{N}}(\mu_{g},\Sigma_{g})$, with probability at least $1-\delta$. In addition, the error in estimating the covariance matrix is bounded by
$\|\widehat{\Sigma}_{g}^{n}-\Sigma_{g}\|_{2}\leq\|\widehat{\mu}_{g}^{n}-\mu_{g}\|_{2}^{2}+\|\Sigma_{g}\|_{2}\left(2\sqrt{\frac{d}{n}}+2\sqrt{\frac{1}{2n}\log\left(\frac{2}{\delta}\right)}+\left(\sqrt{\frac{d}{n}}+\sqrt{\frac{1}{2n}\log\left(\frac{2}{\delta}\right)}\right)^{2}\right)$
which follows from the exact same derivation in Inequality (28) and holds with probability at least $1-\delta$ (see Lemma E.6). Compared to the generator-dependent error bounds utilized by FID-UCB (i.e., Inequalities (27) and (29)), these bounds are dimension-dependent and do not fully exploit the properties of the generator. Finally, by the exact same analysis in the proof of Theorem 5.1, the following estimated FID
$\widehat{\textup{FID}}_{g}^{n}=$ $\widetilde{\textup{FID}}_{g}^{n}-C_{g}^{n}\cdot\sigma_{g}\sqrt{\frac{d}{n}\log\left(\frac{6d}{\delta}\right)}-\textup{Tr}[\Sigma_{g}]\sqrt{\frac{8}{n}\log\left(\frac{6d}{\delta}\right)}$ (32)
$-\textup{Tr}[\Sigma_{\textup{r}}^{\frac{1}{2}}]\sqrt{8\left(\|\Sigma_{g}\|_{2}\left(2\sqrt{\frac{d}{n}}+2\sqrt{\frac{1}{2n}\log\left(\frac{6}{\delta}\right)}+\left(\sqrt{\frac{d}{n}}+\sqrt{\frac{1}{2n}\log\left(\frac{6}{\delta}\right)}\right)^{2}\right)+\frac{d\sigma_{g}^{2}}{n}\log\left(\frac{6d}{\delta}\right)\right)}$
lower bounds $\textup{FID}_{g}$ with probability at least $1-\delta$.
4. Implementation details of naive UCB for IS-based evaluation.
In IS-based evaluation, naive UCB uses Hoeffding-type bonus terms. Specifically, the error vector in defining the optimistic marginal class distribution is given by $e_{g}^{n}$, whose $j$-th element is defined as
$e_{g}^{n}[j]:=\sqrt{\frac{1}{2n}\log\left(\frac{4d}{\delta}\right)},$
which ensures that ${\mathcal{E}}(\textup{Clip}_{e^{-1}}(\widetilde{p}_{Y_{g}}^{n},e_{g}^{n}))\geq H(Y_{g})$ with probability at least $1-\frac{\delta}{2}$. In addition, the concentration error of $\widehat{H}^{n}(Y_{g}|X_{g})$ is bounded by
$H_{g}(Y_{g}|X_{g})\geq\widehat{H}^{n}(Y_{g}|X_{g})-C^{\prime}\sqrt{\frac{1}{2n}\log\left(\frac{4}{\delta}\right)}$
which holds with probability at least $1-\frac{\delta}{2}$. Therefore, the estimated IS is given by
$\widehat{\textup{IS}}_{g}^{n}:=\exp\left\{{\mathcal{E}}\left(\textup{Clip}_{e^{-1}}\left(\widetilde{p}_{Y_{g}}^{n},e_{g}^{n}\right)\right)-\widehat{H}^{n}(Y_{g}|X_{g})+C^{\prime}\sqrt{\frac{1}{2n}\log\left(\frac{4}{\delta}\right)}\right\}$ (33)
By the exact same analysis in the proof of Theorem 6.1, it can be shown that $\widehat{\textup{IS}}_{g}^{n}\geq\textup{IS}_{g}$ with probability at least $1-\delta$.
5. Additional implementation details of FID-UCB
To compute the bonus term (7), we need to estimate the following parameters: 1) $\|\mu_{g}-\mu_{\textup{r}}\|_{2}$: we set it to be a constant, e.g., 0.5; 2) ${\mathcal{I}}_{g}$: we estimate this term using the truncated covariance matrix with parameter $\tau$, which is denoted by $\textup{truncate}_{\tau}(\widehat{\Sigma}_{g}^{n})$. The $(i,j)$-th element of the truncated covariance matrix is $\widehat{\Sigma}_{g}^{n}[i,j]$ if $\widehat{\Sigma}_{g}^{n}[i,j]\geq\tau$ and is zero otherwise. In the experiments, we observe that a truncation parameter that is $5\sim 10\%$ of the operator norm $\|\Sigma_{\textup{r}}\|_{2}$, i.e., $0.05\cdot\|\Sigma_{\textup{r}}\|_{2}\leq\tau\leq 0.1\cdot\|\Sigma_{\textup{r}}\|_{2}$, attains satisfying performance; 3) $\|\Sigma_{g}\|_{2}$: we estimate this term by $\max_{j\in[d]}(\widehat{\sigma}_{j}^{n})^{2}+O(\sqrt{\log(1/\delta)/n})$, where $(\widehat{\sigma}_{j}^{n})^{2}$ is the sample variance for the $j$-th entry of the embedding $f(X_{g})$. This avoids the computation for the eigendecomposition of the sample covariance matrix $\widehat{\Sigma}_{g}^{n}$. The intuition is that when $\Sigma_{g}$ is diagonal, the operator norm $\|\Sigma_{g}\|_{2}$ coincides with the largest per-entry variance; 4) $\textup{Tr}[\Sigma_{g}]$: while a direct approach is to estimate it by $\textup{Tr}[\widehat{\Sigma}_{g}^{n}]$, we found that a better performance can be attained by using $O(\textup{Tr}[\textup{truncate}_{\tau}(\widehat{\Sigma}_{g}^{n})\cdot\Sigma_{\textup{r}}])$, where we utilize the fact that $\|\Sigma_{\textup{r}}^{-1}\|_{2}\cdot\textup{Tr}[\Sigma_{g}\Sigma_{\textup{r}}]\geq\textup{Tr}[\Sigma_{g}]\geq\frac{\textup{Tr}[\Sigma_{g}\Sigma_{\textup{r}}]}{\|\Sigma_{\textup{r}}\|_{2}}$.
Appendix E Auxiliary Lemmas
Lemma E.1 (Concentration of Gaussian random variable).
Let $X_{1},\cdots,X_{n}\sim{\mathcal{N}}(\mu,\sigma^{2})$. Then, it holds that
${\mathbb{P}}\left(\left|\frac{1}{n}\sum_{i=1}^{n}X_{i}-\mu\right|>\epsilon\right)<\exp\left(-\frac{n\epsilon^{2}}{2\sigma^{2}}\right)$
Equivalently, with probability at least $1-\delta$, we have that
$\left|\frac{1}{n}\sum_{i=1}^{n}X_{i}-\mu\right|\leq\sqrt{\frac{\sigma^{2}}{2n}\log\left(\frac{2}{\delta}\right)}.$
Lemma E.2 (Hoeffding’s inequality).
Let $Z,Z_{1},\cdots,Z_{n}$ be i.i.d. random variables with values in $[0,1]$ and let $\delta>0$. Then with probability at least $1-\delta$ in $(Z_{1},\cdots,Z_{n})$ we have that
$\mathbb{E}Z-\frac{1}{n}\sum_{i=1}^{n}Z_{i}\leq\sqrt{\frac{1}{2n}\log\left(\frac{1}{\delta}\right)}.$ (34)
A drawback of Hoeffding’s inequality is that the confidence interval is distribution-agnostic, and always scales with order $O(\sqrt{1/n})$. For distribution with small variance, the following [18, Bennett’s inequality] provides a tighter concentration error.
Lemma E.3 (Bennett’s inequality).
Under the conditions of Theorem E.2, with probability at least $1-\delta$ in $(Z_{1},\cdots,Z_{n})$ we have that
$\mathbb{E}Z-\frac{1}{n}\sum_{i=1}^{n}Z_{i}\leq\sqrt{\frac{2{\mathbb{V}}Z}{n}\log\left(\frac{1}{\delta}\right)}+\frac{1}{3n}\log\left(\frac{1}{\delta}\right),$ (35)
where ${\mathbb{V}}Z$ is the variance ${\mathbb{V}}=\mathbb{E}(Z-\mathbb{E}Z)^{2}$.
When the exact variance is unknown, [31] provide a purely data-dependent bound with similar properties as Bennett’s inequality.
Lemma E.4 (Restatement of [31, Theorem 4]).
Under the conditions of Theorem E.2, we have that with probability at least $1-\delta$ in the i.i.d vector $\bm{Z}=(Z_{1},\cdots,Z_{n})$ that
$\mathbb{E}Z-\frac{1}{n}\sum_{i=1}^{n}Z_{i}\leq\sqrt{\frac{2V_{n}(\bm{Z})}{n}\log\left(\frac{2}{\delta}\right)}+\frac{7}{3(n-1)}\log\left(\frac{2}{\delta}\right),$
where $V_{n}(\bm{Z}):=\frac{1}{n(n-1)}\sum_{1\leq i<j\leq n}(Z_{i}-Z_{j})^{2}$ is the sample variance.
Lemma E.5 (Concentration of Chi-square variables).
Let $Z_{1},\cdots,Z_{n}\sim{\mathcal{N}}(0,1)$ be i.i.d standard normal. Then, it holds that
${\mathbb{P}}\left(\left|\frac{1}{n}\sum_{i=1}^{n}Z_{i}^{2}-1\right|>\epsilon\right)\leq 2e^{-n\epsilon^{2}/8}.$
Equivalently, with probability at least $1-\delta$, we have that
$\left|\frac{1}{n}\sum_{i=1}^{n}Z_{i}^{2}-1\right|\leq\sqrt{\frac{8}{n}\log\left(\frac{2}{\delta}\right)}$
Lemma E.6 (Restatement of [43, Example 6.3]).
Let $\bm{X}\in{\mathbb{R}}^{n\times d}$ be a random matrix from the $\Sigma$-Gaussian ensemble, i.e., each row vector $x_{i}$ is drawn i.i.d. from a multivariate ${\mathcal{N}}(0,\Sigma)$ distribution. Then, with probability at least $1-\delta$, we have that
$\|\widehat{\Sigma}-\Sigma\|_{2}\leq\|\Sigma\|_{2}\left(2\sqrt{\frac{d}{n}}+2\epsilon+\left(\sqrt{\frac{d}{n}}+\epsilon\right)^{2}\right),\textup{ where }\epsilon=\sqrt{\frac{1}{2n}\log\left(\frac{2}{\delta}\right)},$ (36)
where $\widehat{\Sigma}=\frac{1}{n}\sum_{i=1}^{n}x_{i}x_{i}^{\top}$ is the sample covariance matrix.
Lemma E.7 (Dimension-free concentration of sample covariance matrix [30, 25, 46]).
Let $X_{1},\cdots,X_{n}$ are i.i.d zero-mean and sub-Gaussian random vectors with parameter at most $\sigma^{2}$, whose covariance matrix is given by $\Sigma$. Let
$\bm{r}(\Sigma):=\frac{\textup{Tr}(\Sigma)}{\|\Sigma\|_{2}}$ (37)
denote the effective rank. Then, with probability at least $1-\delta$, the sample covariance matrix satisfies that
$\left\|\frac{1}{n}\sum_{i=1}^{n}X_{i}X_{i}^{\top}-\Sigma\right\|_{2}\leq 20\sigma^{2}\|\Sigma\|_{2}\sqrt{\frac{4\bm{r}(\Sigma)+\log(1/\delta)}{n}}$ (38)
whenever $n\geq 4\bm{r}(\Sigma)+\log(1/\delta)$.
Lemma E.8.
Let $A,B\in{\mathbb{S}}^{d}$ be positive semi-definite (PSD) matrices. Then, we have that
$\textup{Tr}(\sqrt{A+B})\geq\max\{\textup{Tr}(\sqrt{A}),\textup{Tr}(\sqrt{B})\}$ (39)
Proof.
Since $A+B\succeq A$ and $A+B\succeq B$, by Courant-Fischer min-max theorem, it holds that $\lambda_{i}(A+B)\geq\lambda_{i}(A)$ and $\lambda_{i}(A+B)\geq\lambda_{i}(B)$ for their $i$-th largest eigenvalues, which concludes the proof. ∎
Lemma E.9.
Let $A,B\in{\mathbb{S}}^{d}$ be positive semi-definite (PSD) matrices. Then, we have that
$\textup{Tr}[\sqrt{A+B}]\leq\textup{Tr}[\sqrt{A}]+\textup{Tr}[\sqrt{B}].$ (40)
Proof.
First, we assume that both matrices $A$ and $B$ are positive strictly definite. Then,
$\textup{Tr}(\sqrt{A+B})=$ $\textup{Tr}[(A+B)^{-\frac{1}{4}}(A+B)(A+B)^{-\frac{1}{4}}]$
$=$ $\textup{Tr}[(A+B)^{-\frac{1}{4}}A(A+B)^{-\frac{1}{4}}]+\textup{Tr}[(A+B)^{-\frac{1}{4}}B(A+B)^{-\frac{1}{4}}]$
$=$ $\textup{Tr}[\sqrt{A}(A+B)^{-\frac{1}{2}}\sqrt{A}]+\textup{Tr}[\sqrt{B}(A+B)^{-\frac{1}{2}}\sqrt{B}]$
$\leq$ $\textup{Tr}[\sqrt{A}]+\textup{Tr}[\sqrt{B}],$
where the last inequality holds by $A^{-\frac{1}{2}}\succeq(A+B)^{-\frac{1}{2}}$ and $B^{-\frac{1}{2}}\succeq(A+B)^{-\frac{1}{2}}$ (see Lemma E.12). For general PSD matrices $A$ and $B$, note that
$\textup{Tr}(\sqrt{A+B+\epsilon I})\leq\textup{Tr}(\sqrt{A+\epsilon I})+\textup{Tr}(\sqrt{B+\epsilon I})$
holds for any $\epsilon>0$ by our previous analysis. Since the trace map is continuous, ∎
Lemma E.10.
Let $z_{1},\cdots,z_{n}\sim{\mathcal{N}}(0,1)$ are i.i.d standard Gaussian. Then, with probability at least $1-\delta$, it holds that
$\left(\frac{1}{n}\sum_{i=1}^{n}z_{i}\right)^{2}\leq\sqrt{\frac{32}{n^{2}}\log\left(\frac{2}{\delta}\right)}.$
Proof.
Note that $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}z_{i}\sim{\mathcal{N}}(0,1)$. By Lemma E.5, we have that
${\mathbb{P}}\left(\left|\left(\frac{1}{n}\sum_{i=1}^{n}z_{i}\right)^{2}-\frac{1}{n}\right|>\epsilon\right)=$ ${\mathbb{P}}\left(|U^{2}-1|>n\epsilon\right)\leq 2\exp\left(-\frac{n^{2}\epsilon^{2}}{8}\right)$
where $U^{2}\sim\chi^{2}_{1}$ is a chi-square random variable. Hence, with probability at least $1-\delta$, it holds that
$\left|\left(\frac{1}{n}\sum_{i=1}^{n}z_{i}\right)^{2}-\frac{1}{n}\right|\leq\sqrt{\frac{8}{n^{2}}\log\left(\frac{2}{\delta}\right)},$
which concludes the proof. ∎
Lemma E.11.
Let $z_{1},\cdots,z_{n},u_{1},\cdots,u_{n}\sim{\mathcal{N}}(0,1)$ be i.i.d standard Gaussian. Then, with probability at least $1-\delta$, it holds that
$\left|\left(\frac{\sum_{i=1}^{n}z_{i}}{n}\right)\left(\frac{\sum_{i=1}^{n}u_{i}}{n}\right)\right|<\frac{1}{n}\sqrt{8\log\left(\frac{4}{\delta}\right)}.$
Proof.
Note that $\frac{1}{\sqrt{n}}\sum_{i=1}^{n}z_{i}\sim{\mathcal{N}}(0,1)$.
$\left(\frac{\sum_{i=1}^{n}z_{i}}{n}\right)\left(\frac{\sum_{i=1}^{n}u_{i}}{n}\right)=\frac{1}{n}\left(\frac{\sum_{i=1}^{n}z_{i}}{\sqrt{n}}\right)\left(\frac{\sum_{i=1}^{n}u_{i}}{\sqrt{n}}\right)=\frac{1}{n}ZU,$
where $Z,U\sim{\mathcal{N}}(0,1)$ are i.i.d standard Gaussian. Observe that
$ZU=\frac{1}{4}(Z+U)^{2}-\frac{1}{4}(Z-U)^{2}\sim\frac{1}{2}Q-\frac{1}{2}R,$
where $Q:=\frac{1}{2}(Z+U)^{2},R:=\frac{1}{2}(Z-U)^{2}\sim\chi_{1}^{2}$ are i.i.d Chi-square distributed with 1 degree of freedom (since $\frac{1}{\sqrt{2}}(Z+U)$ and $\frac{1}{\sqrt{2}}(Z-U)$ are i.i.d standard Gaussian). Hence, by Lemma E.5 and a union bound, with probability at least $1-\delta$,
$\left|\left(\frac{\sum_{i=1}^{n}z_{i}}{n}\right)\left(\frac{\sum_{i=1}^{n}u_{i}}{n}\right)\right|=$ $\left|\frac{1}{2}Q-\frac{1}{2}R\right|\leq\frac{1}{2}\left|Q-1\right|+\frac{1}{2}\left|R-1\right|$
$\leq$ $\frac{1}{n}\sqrt{8\log\left(\frac{4}{\delta}\right)},$
which concludes the proof. ∎
Lemma E.12.
Let $A,B$ denote two PSD matrices such that $A\succeq B$. Then, their inverse $B^{-1}\succeq A^{-1}$.
Proof.
It suffices to show that $\sqrt{A}B^{-1}\sqrt{A}\succeq I$. Since $\sqrt{A}B^{-1}\sqrt{A}=(\sqrt{A}B^{-\frac{1}{2}})(\sqrt{A}B^{-\frac{1}{2}})^{\top}$ shares the same eigenvalues with $(\sqrt{A}B^{-\frac{1}{2}})^{\top}(\sqrt{A}B^{-\frac{1}{2}})=B^{-\frac{1}{2}}AB^{-\frac{1}{2}}\succeq I$, which concludes the proof. ∎